# ROVR-Open-Dataset: A Large-Scale Depth Dataset for Autonomous Driving

### v1

@@ IV-C Main Results
As shown in Tables III, models trained on KITTI [10], DDAD [42], or nuScenes [8] achieve strong results within their own domains, yet when transferred to ROVR, their performance drops dramatically across all metrics. Table IV reveal the saturation of existing benchmarks. SOTA methods that excel on KITTI [10] benchmarks still struggle on ROVR, showing that current approaches cannot fully adapt to its diverse and challenging scenarios. Training on ROVR yields consistent improvements across all evaluation metrics on ROVR, underscoring the critical role of scale and diversity in enabling models to generalize effectively to complex real-world scenarios. However, the absolute performance of existing methods remains unsatisfactory — for instance, the best abs_rel is still around 0.2 — indicating that current depth estimation approaches have already reached saturation on traditional datasets. This highlights the necessity of developing more practical models that can effectively learn from sparse depth ground truth and better handle the challenges presented by ROVR.
@@ IV-D Ablation Study
Table V presents depth estimation performance across different illumination conditions and scene types, demonstrating the ROVR dataset’s diversity and the varying challenges posed by different environments.
@@
Results under varying illumination conditions. The upper block of the table shows results under varying illumination conditions—Normal, Night, and Rainy. As expected, all methods perform better in normal lighting conditions, where visibility and depth prediction quality are optimal. However, the performance significantly declines in low-light conditions, especially at night, as seen in the reduction of metrics like $\delta_{1}$ and $\delta_{2}$. VA-DepthNet consistently outperforms other methods in night-time settings, achieving the lowest abs_rel and rms values. The Rainy condition shows the largest performance drop, with all methods facing challenges due to degraded RGB quality and increased LiDAR sparsity. In rainy conditions, IEBins struggles the most, with substantially higher log10 and sq_rel values, indicating poor performance in handling noise and sparse data.
@@
Results across different scene types. The lower block of the table reports results across different scene types—Highway, Rural, and Urban. Highway scenes, characterized by relatively clear and structured environments, yield the best results for all methods, with VA-DepthNet leading the pack in terms of both $\delta_{1}$ and silog scores. This can be attributed to the simpler geometry and better lighting consistency typically found on highways. Conversely, Urban scenes present the greatest difficulty for depth estimation. With complex urban layouts featuring dense clutter, frequent occlusions, and multi-scale objects, IEBins shows the largest performance gap, with noticeably higher error values across all metrics. These results highlight the challenges posed by dynamic and heterogeneous urban environments, where traditional methods struggle to maintain accuracy.
Influence of the Scale. Fig. 3 demonstrates that as the number of test identities increases from 2K to 10K, the performance curves remain relatively stable. This indicates that the ROVR dataset provides a consistent and unbiased evaluation environment, where scaling the test set does not distort the overall difficulty. Such stability highlights the dataset’s reliability as a robust benchmark for assessing depth estimation methods at large scale.
@@ IV-E Qualitative Results
Fig. 4 showcases qualitative samples from the ROVR dataset across multiple scene types (highway, rural, urban) and varying illumination and weather conditions (normal, night, rainy). Examples are drawn from different illumination conditions and scene types to match the settings in Table V. These examples highlight the dataset’s challenging nature: adverse weather and nighttime conditions significantly degrade image quality and reduce ground-truth density, while urban environments introduce dense clutter, occlusions, and multi-scale structures. Such diversity and difficulty expose clear failure modes in existing models, indicating that ROVR not only reflects real-world driving challenges but also sets a demanding benchmark for the next generation of depth estimation methods.

### v2

@@ IV-C Main Results
As shown in Tables III, models trained on KITTI [10], DDAD [49], or nuScenes [8] achieve strong results within their respective domains but suffer severe degradation when evaluated on ROVR, revealing poor cross-dataset generalization. In particular, metrics such as abs_rel increase by an order of magnitude, and accuracy drops drastically (e.g., $\delta_{1}$ falls from >0.98>0.98 on KITTI to <0.25<0.25 on ROVR).
These results expose two key issues. First, existing benchmarks are already saturated—state-of-the-art methods achieve near-ceiling scores on KITTI—yet the same models fail to generalize to ROVR, demonstrating limited robustness. Second, even when trained directly on ROVR, current methods perform far from saturating the benchmark (best $\text{abs\_rel}\sim 0.18$–0.21), underscoring that ROVR presents a substantially more challenging setting. This establishes ROVR as a more realistic and difficult benchmark, motivating the need for new depth estimation approaches capable of handling its scale, diversity, and sparse ground truth.
@@ IV-D Ablation Study
Table V and Table VI present depth estimation performance across different illumination conditions and scene types. We report the results of three typical depth estimation methods to provide a more complete understanding.
Results under varying illumination conditions. Table V shows results under varying illumination conditions—Normal, Night, and Rainy. As expected, all methods perform better in normal lighting conditions, where visibility and depth prediction quality are optimal. However, the performance significantly declines in low-light conditions, especially at night, as seen in the reduction of metrics like $\delta_{1}$ and $\delta_{2}$. VA-DepthNet consistently outperforms other methods in night-time settings, achieving the lowest abs_rel and rms values. The Rainy condition shows the largest performance drop, with all methods facing challenges due to degraded RGB quality and increased LiDAR sparsity. In rainy conditions, IEBins struggles the most, with substantially higher log10 and sq_rel values, indicating poor performance in handling noise and sparse data.
Results across different scene types. Table VI reports results across different scene types—Highway, Rural, and Urban. Highway scenes, characterized by relatively clear and structured environments, yield the best results for all methods, with VA-DepthNet leading the pack in terms of both $\delta_{1}$ and silog scores. This can be attributed to the simpler geometry and better lighting consistency typically found on highways. Conversely, Urban scenes present the greatest difficulty for depth estimation. With complex urban layouts featuring dense clutter, frequent occlusions, and multi-scale objects, IEBins shows the largest performance gap, with noticeably higher error values across all metrics. These results highlight the challenges posed by dynamic and heterogeneous urban environments, where traditional methods struggle to maintain accuracy.
@@
Influence of the Test Scale. Fig. 3 demonstrates that as the number of test identities increases from 2K to 10K, the performance curves remain relatively stable. This indicates that the ROVR dataset provides a consistent and unbiased evaluation environment, where scaling the test set does not distort the overall difficulty. Such stability highlights the dataset’s reliability as a robust benchmark for assessing depth estimation methods at a large scale.
@@ IV-E Qualitative Results
Fig. 4 showcases qualitative samples from the ROVR dataset across multiple scene types (highway, rural, urban) and varying illumination and weather conditions (normal, night, rainy). Examples are drawn from different illumination conditions and scene types to match the settings in Table V and Table VI. These examples highlight the dataset’s challenging nature: adverse weather and nighttime conditions significantly degrade image quality and reduce ground-truth density, while urban environments introduce dense clutter, occlusions, and multi-scale structures. Such diversity and difficulty expose clear failure modes in existing models, indicating that ROVR not only reflects real-world driving challenges but also sets a demanding benchmark for the next generation of depth estimation methods.